### v1

# Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-tuning

Abstract
Recent vision-language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images. However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns. While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage” where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm. We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards. Additionally, these correlations contribute to the over prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily. To address this issue, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities. Extensive evaluations across safety benchmarks show that under one-word attacks, MU-based alignment reduces the attack success rate by up to 60.17% and cuts unnecessary rejections by over 84.20%. Codes are available at https://github.com/OPTML-Group/VLM-Safety-MU. WARNING: There exist AI generations that may be offensive in nature.
1 Introduction
Recent multi-modal models have achieved advancements by integrating text and images [1, 2, 3, 4, 5]. A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module [6, 7, 8, 9, 10]. While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities [11, 12, 13, 14, 15]. Specifically, images can function as a “foreign language” [16], creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety [16, 17, 18]. The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), provided that a high-quality, dual-modality curated safety fine-tuning dataset is available [14, 19, 20, 21, 22]. One compelling piece of evidence is that fine-tuning on VLGuard [23], a widely used VLM safety dataset, substantially improves robustness against unsafe queries and jailbreaking attacks. As demonstrated in [23], this enhancement surpasses the results obtained by fine-tuning on a “clean” dataset which removes the unsafe data. The surprising effectiveness of SFT on VLGuard has sparked growing interest in re-evaluating its reliability. Recent studies [18, 11, 24] have identified a downside of such safety fine-tuning, known as the over-prudence problem. This issue refers to the over-conservatism of VLMs after safety fine-tuning, where they unnecessarily reject responses when presented with benign queries.
However, the over-prudence suggests that these models may be overly safe, to avoid responses to any skeptical queries. Alternatively, the observed safety may be illusory, as current safety fine-tuning fails to ensure reliability, giving a false sense of safety. Thus, we ask:
(Q) Does current VLM safety fine-tuning achieve true safety? If not, what is the root cause?
Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed, fine-tuned on VLG uard [23]) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in the original unsafe query) can bypass the safety mechanism established through fine-tuning on VLGuard , even though the safeguarded VLM correctly rejects the original unsafe query. (b) Over-prudence issue: Similar to the one-word attack, a minor modification by replacing “What” with “Share” can cause unnecessary refusals even for benign queries. (c) Root cause of spurious correlations given by fine-tuning dataset biases: Certain words become disproportionately associated with specific safety labels. For example, “Share” is strongly correlated with rejection responses, while “What” is highly associated with non-rejection responses. (d) Effectiveness of unlearning-based safety fine-tuning: The unlearning methods NPO [25] and RMU [26] enhance robustness against attacks while reducing over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed [23] (“Prior work”).
In this work, we investigate (Q) and challenge the prevailing belief in the effectiveness of safety fine-tuning for VLMs. We uncover a “safety mirage” in VLM safety fine-tuning, where the seemingly robust safety performance after fine-tuning is primarily driven by spurious correlations between certain non-essential textual words in input queries and predefined safety labels (e.g., rejection) in the fine-tuning dataset. If an adversary identifies these spurious correlations, a simple one-word modification–which we refer to as the “one-word attack”–can effectively jailbreak safety fine-tuned VLMs, enabling them to regenerate unsafe content. Additionally, the input-rejection label shortcut induced by these spurious correlations provides an explanation for the over-prudence of safety fine-tuned VLMs. Similar to the one-word attack, a one-word modification in text queries can readily activate the input-rejection shortcut at test time, causing the model to overgeneralize rejection responses and refuse to generate outputs even for benign queries. In Fig. 1(a)-(c), we provide a schematic overview illustrating: (a) The one-word attack on a safety fine-tuned VLM LLaVA-v1.5-7B-Mixed [23]; (b) The over-prudence issue by one-word modification; (c) The spurious correlations observed in the fine-tuning dataset VLGuard, where the word “share” is strongly linked to rejection, while “what” is associated with non-rejection.
Building on our identification of spurious correlations in VLM safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU). Originally designed to remove the influence of undesired data or knowledge from ML models, MU ensures that essential knowledge remains intact while avoiding unintended disruptions to causally unrelated information [27, 28, 29]. We propose adapting MU to VLM safety fine-tuning as a more robust alternative to traditional supervised approaches. Rather than enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels. Although MU has been applied to VLM safety fine-tuning in prior work [17, 30, 31], its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored. Fig. 1(d) showcases the effectiveness of applying two LLM unlearning approaches, NPO [25] and RMU [26] in enhanced robustness against jailbreaking attacks and reduced over-prudence rates.
In summary, our key contributions are listed below.
① We revisit the problem of supervised safety fine-tuning for VLMs and find that there exists a safety mirage, driven by hidden biases—specifically, spurious correlations between non-essential textual features and safety labels in the fine-tuning dataset.
② From an attack perspective, we show that safety fine-tuned VLMs are still susceptible to jailbreaking when adversaries exploit spurious correlations embedded in the fine-tuning dataset. We propose a simple and effective one-word attack by substituting highly frequent querying words associated with rejection responses with those linked to normal model outputs. Additionally, we show that these spurious correlations also contribute to over-prudence, causing fine-tuned VLMs to unnecessarily reject benign inputs.
③ From a defense perspective, we show that MU offers a promising solution to alleviate the effects of spurious correlations in fine-tuning data for VLM safety fine-tuning. The key rationale is that unlearning removes the influence of unsafe responses without relying on the spurious feature-label correlation present in the fine-tuning dataset.
④ We conduct extensive experiments across multiple VLM safety evaluation benchmarks, including VLGuard [23], SPA-VL [21], MM-SafetyBench [14], and FigStep [15], and assess model utility on standard VQA datasets. Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
2 Related Work
VLM safety: Attack and defense. With the rapid advancement of VLMs [6, 7, 8, 32, 33, 34, 1, 2, 4], safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content. While LLMs have been extensively studied for safety risks, leading to the development of attack strategies [35, 36, 37, 38], defense mechanisms [39, 40, 41], and robust evaluation datasets [42, 43, 44], VLMs introduce additional challenges due to the complexity of multimodal inputs [16, 17, 18], making them even more vulnerable to jailbreaking and adversarial manipulation [11, 12, 13, 14, 15]. Attacks on VLMs often leverage the dual-modality nature of these models. One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs [15, 14]. Another strategy involves using gradient-based adversarial image generation [45, 46, 47, 48, 49] to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning. Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets. The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms [50, 51, 16, 52, 24]. The latter focuses on red-teaming dataset curation [14, 19, 20, 21, 22, 23, 53], enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
Machine unlearning in VLMs. MU [27, 28, 29, 54, 55] is designed to remove harmful data influences from a pre-trained model while preserving its overall utility. In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance [25, 26, 56, 57, 58, 59, 60]. In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms [61, 62, 63]. For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation [30, 31, 17]. Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
3 Preliminary and Problem Statement
Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc [23, 64]). Green shield represents the correct response, either a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries. And red question ? represents an inappropriate rejection for a safe query. (a) Successfully jailbreaking: The safety fine-tuned model originally produces rejection-based responses for unsafe queries; However, replacing the initial question word with “What” can easily bypass this safeguard. (b) Over-prudence: A minor modification by replacing “What” with “Share” can trigger unnecessary refusals even for benign queries.
Existing VLM safety fine-tuning setup.
Previous works [16, 15, 14] highlight the need for safety alignment in VLMs to encompass both textual and imagery data. Consequently, many efforts [23, 21, 65] have focused on curating high-quality dual-modality safety datasets for VLMs. A notable benchmark dataset is VLGuard [23], which covers various text-image pairing scenarios, including unsafe cases where either the text is unsafe or both text and image are unsafe, as well as safe cases where both modalities are benign. Leveraging these curated safety datasets, recent works [23, 17, 18, 21] show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks [66, 67, 68]. In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
We begin by presenting the problem formulation of VLM safety fine-tuning. Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard). In addition, let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, which consists of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning. For a VLM parameterized by ${\bm{\theta}}$, the safety fine-tuning problem can be formulated as:
$\begin{array}[]{ll}\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}&\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})+\gamma\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),\end{array}$ (2)
where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and the regularization parameter $\gamma\geq 0$ strikes a balance between safety alignment and preserving performance on normal tasks.
Safety mirage: Motivation and problem of interest. Although the safety fine-tuned VLMs could overly reject queries even when presented with benign ones [18, 11, 24], the resulting safety performance against unsafe queries does not degrade; It remains highly robust even in the presence of some common jailbreaking attacks [23, 21].
The seemingly ‘robust’ safety performance observed after fine-tuning motivates us to re-examine its true reliability. As demonstrated in Fig. 1, the current safety fine-tuned VLM remains highly vulnerable to simple paraphrasing of text queries even when only the first question word is modified. As a supplement to Fig. 1, Fig. 2 provides additional motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence. Consistent with Fig. 1, unsafe queries prefixed with innocuous question word “What” successfully bypass safeguard, and harmless prompts start with “Share” trigger over-rejection effect. Notably, the choice of the replacement word ‘What’ or ‘Share’ is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec. 4.
Examples in Figs. 1-2 suggest that fine-tuning VLMs on safety datasets may create a “safety mirage”, as evidenced by their susceptibility to even minor one-word modification in text queries. Thus, our work focuses on the following key research questions: (a) What is the root cause of the “safety mirage” in VLM safety fine-tuning? (b) What can be improved to mitigate the “safety mirage”? We explore these questions in Sec. 4 and Sec. 5, respectively.
4 The Risks of Spurious Correlations in VLM Safety Fine-Tuning
Spurious text features and spurious correlations. As discussed in Sec. 3, the effectiveness of current VLM safety fine-tuning methods heavily relies on the curation of high-quality, dual-modality safety datasets. As a result, the safety capabilities of fine-tuned VLMs (i.e., their abilities to prevent harmful content generation) are primarily learned from safety labels (i.e., safe responses) introduced in the fine-tuning datasets. For example, in VLGuard [23], safety labels for unsafe text queries and/or unsafe images are assigned as rejection responses, such as “I’m sorry, I cannot assist with that request…” against the unsafe text query in Figs. 1-2. Additionally, safety labels may correspond to standard VLM responses when processing safe text-image inputs.
At first glance, the use of safety labels appears appropriate. However, a hidden bias may arise when these safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work. Here the term “spurious features” refers to non-essential textual words in inputs that do not contribute to the fundamental meaning or task-relevant aspects of the input query. For example, in Fig. 2, the word “What” or “Share” at the beginning of the input query can be considered a spurious feature because it does not directly relate to the query’s actual content and can be easily substituted with other question words. In contrast, “core” features (such as “crime”) are more informative, representing content-related words that capture the true intent and meaning of the query. Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the assigned labels in the safety fine-tuning dataset.
The above conceptualization of spurious correlations is inspired by conventional analyses in image classification [69, 70, 71], where spurious features correspond to background or non-object pixels, while core features directly relate to the target object. In these studies, spurious correlations arise when the target label is linked to spurious features, rather than solely to core features. In the context of VLMs, while several studies have explored spurious correlations [72, 73], their focus has been primarily on spurious visual features within the image modality. To the best of our knowledge, the spurious correlations caused by textual features in VLM safety fine-tuning have not been investigated before.
In this work, we identify two types of spurious correlations in VLM safety fine-tuning. (a) Non-rejection bias: Certain words (like “What” in Fig. 2(a)) in text queries become spuriously correlated with non-rejection responses. As a result, incorporating these words into an original query can easily jailbreak fine-tuned VLMs. (b) Rejection bias: Certain words (like “Share” in Fig. 2(b)) in text queries become spuriously correlated with rejection responses, causing the fine-tuned VLM to exhibit over-prudence.
Refer to caption Refer to caption
(a) Top words in safe queries (b) Top words in unsafe queries
Figure 3: Frequency of question-initiating words in VLGuard queries: (a) words in safe queries associated with non-rejection responses, (b) words in unsafe queries associated with rejections.
To determine the spurious textual features and the spurious correlations (a)-(b), we analyze the frequency of words used in train-time text queries that lead to non-rejection responses (i.e., generating textual content in response to safe queries in the training set) and rejection responses (i.e., predefined refusal answers against unsafe queries in the training set), respectively. Fig. 3 presents the most frequently occurring starting words in text queries from the VLGuard dataset, categorized based on their tendency to elicit non-rejection responses or rejection responses. As shown, the question word “what” predominantly correlates with non-rejection responses, appearing in over 80% of safe queries where the model generates a response. In contrast, the question-initiating words “can” and “share” in unsafe queries are strongly associated with rejection responses, with over 50% of its occurrences leading to a rejection. Compared to “can”, which appears in both safe and unsafe queries, the word “share” is only used in unsafe queries to elicit rejection.
One-word jailbreaking. Recognizing the non-rejection bias (i.e., the spurious correlation between certain querying words like “what” and non-rejection responses) as shown in Fig. 3, an adversary can exploit this spurious correlation to jailbreak safety fine-tuned VLMs. Formally, let $q$ denote the original unsafe text query that VLM avoids generating unsafe content for. And let $q^{\prime}=\mathrm{w}_{\mathrm{adv}}+q$ denote the jailbreaking attack, where $\mathrm{w}_{\mathrm{adv}}$ is the adversarial perturbation chosen based on the non-rejection bias-inducing querying word (e.g., “what” in Figs. 1, 2, and 3), and the operation $+$ signifies either word insertion as a prefix to $q$ or a simple starting word replacement in $q$. We refer to the above simple attack strategy as the one-word jailbreaking attack.
Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed [23]).
[…]
One-word over-prudence. Similar to the one-word jailbreaking attack, when a rejection bias-inducing word (like “Share” in Fig. 3) is introduced into text queries, we observe that even benign modified queries can still prevent the fine-tuned VLM from generating any meaningful output. To amplify the over-prudence rate, we can also apply the multi-shot strategy (i.e., “Share”-based one-word modification to different paraphrased versions of the benign input query). The results for the multi-shot setting are presented in Appendix A. Notably, even in the one-shot setting, the “Share”-initialized benign query modification already achieves a 90% over-rejection rate against safe text-image queries.
5 Enhancing VLM Safety through Unlearning
To mitigate spurious correlations embedded in the safety fine-tuning dataset, one potential solution, motivated by the analysis in Sec. 4, is to eliminate the dependence of safety fine-tuning on safety labels (e.g., rejection responses to unsafe queries). This necessitates shifting from supervised fine-tuning to a label-free, unsupervised setting for safety alignment. Machine unlearning (MU) [27, 28, 29] provides an ideal solution in this context, as it is designed to remove the undesired influence of harmful data or knowledge from a pre-trained model while preserving its normal utility.
Although unlearning has been applied to VLM safety fine-tuning in prior work [17, 30, 31], its unique advantage and application in circumventing spurious correlations within fine-tuning datasets remains unexplored. To this end, we adapt two state-of-the-art MU approaches, originally developed for large language models (LLMs)–representation misdirection unlearning (RMU) [26] and negative preference optimization (NPO) [25]–to the context of VLM unlearning.
The proposed VLM unlearning follows the generic formulation of (2), but with the following key modifications. First, the fine-tuning loss over the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ is replaced with an unlearning objective $\ell_{\mathrm{u}}$ that relies solely on the unsafe data features (text-image queries) in $\mathcal{D}_{\mathrm{u}}$, without depending on the safety labels. In our work, we define the unlearning loss $\ell_{\mathrm{u}}$ based on the principles of RMU and NPO, respectively. The RMU-based unlearning objective aims to map the intermediate features of unsafe data $x\in\mathcal{D}_{\mathrm{u}}$ (to be forgotten) to random features. This ensures that the model no longer retains meaningful representations of the unsafe data. The objective is given by
$\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}[\|M_{{\bm{\theta}}}(\mathbf{x})-c\cdot\mathbf{v}\|_{2}^{2}],$ (3)
where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of ${\bm{\theta}}$, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. We remark that, unlike RMU for LLM unlearning [26], we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
Table 1: Experiment results evaluating safety, over-prudence, and utility of safety fine-tuned VLMs. Safety is quantified by ASR (attack success rate) on unsafe input queries, evaluated before and after a 3-shot one-word attack (i.e., “what”-based prefix in Fig. 1) that promotes non-rejection bias. Over-prudence is measured by RR (rejection rate) on safe input queries, evaluated before and after using 1-shot, one-word over-prudence modification (i.e., “share”-based prefix in Fig. 1). Here, “Before” and “After” denote the performance prior to and following the respective one-word modification. Utility is assessed by the accuracy (Acc.) on four VQA benchmarks: VQAv2, TextVQA, ScienceQA, and VizWiz. Results are presented for models under full fine-tuning and LoRA fine-tuning settings, with safety fine-tuning approaches including Unsafe-Filter, Mixed-SFT, Posthoc-SFT, NPO-Unlearning, and RMU-Unlearning.
Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence Evaluation (RR, $\downarrow$) Utility Evaluation (Acc., $\uparrow$)
VLGuard SPA-VL VLGuard SPA-VL VQAv2 TextVQA ScienceQA VizWiz
Before After Before After Before After Before After
LLaVA-1.5-7B 64.25% 90.27% 46.42% 52.08% 0.36% 0.36% 14.72% 9.81% 78.53% 58.23% 69.51% 50.07%
       + Unsafe-Filter 65.66% 90.72% 45.66% 54.72% 0.36% 0.36% 15.85% 11.32% 79.14% 58.22% 68.12% 52.14%
       + Mixed-SFT 0.23% 54.98% 14.34% 37.73% 4.48% 91.76% 68.68% 98.87% 78.23% 57.80% 68.27% 52.94%
       + Posthoc-SFT 0.23% 46.83% 13.58% 32.96% 2.69% 90.83% 60.38% 100.0% 78.03% 57.73% 68.42% 51.84%
       + NPO-Unlearning 2.49% 12.92% 18.49% 24.15% 2.51% 11.69% 16.60% 17.36% 77.34% 57.80% 68.02% 50.21%
       + RMU-Unlearning 1.29% 10.18% 17.73% 22.64% 1.25% 7.56% 18.11% 19.24% 77.04% 56.89% 67.68% 50.01%
LLaVA-1.5-7B-LoRA 64.72% 95.25% 44.91% 50.44% 0.18% 0.18% 15.47% 12.45% 79.13% 58.22% 68.62% 52.82%
       + Unsafe-Filter 67.19% 93.89% 45.28% 52.33% 0.36% 0.0% 22.64% 13.21% 79.14% 57.66% 67.97% 53.65%
       + Mixed-SFT 0.45% 69.23% 21.51% 40.13% 3.05% 89.93% 59.25% 97.36% 78.63% 57.24% 68.47% 51.84%
       + Posthoc-SFT 0.23% 51.81% 20.38% 37.61% 3.41% 95.14% 62.26% 99.62% 78.23% 57.17% 67.92% 52.08%
       + NPO-Unlearning 4.56% 18.29% 21.51% 25.28% 2.69% 11.01% 16.98% 19.62% 77.32% 56.98% 66.98% 51.01%
       + RMU-Unlearning 3.87% 11.14% 20.38% 24.24% 1.25% 4.84% 18.49% 21.89% 76.99% 56.62% 66.32% 49.87%
LLaVA-1.5-13B 68.10% 91.86% 50.19% 54.47% 0.54% 0.72% 19.62% 14.34% 79.99% 61.25% 72.73% 53.64%
       + Unsafe-Filter 67.65% 92.99% 52.08% 56.27% 0.54% 0.54% 20.38% 15.09% 79.87% 61.32% 71.59% 52.68%
       + Mixed-SFT 0.45% 57.01% 18.11% 40.5% 4.84% 92.63% 58.87% 97.74% 79.03% 60.98% 72.03% 53.01%
       + Posthoc-SFT 1.58% 69.23% 16.98% 32.83% 2.69% 76.08% 56.98% 98.49% 78.94% 60.63% 71.94% 52.31%
       + NPO-Unlearning 1.89% 11.70% 22.26% 26.04% 2.33% 10.65% 23.77% 27.92% 78.31% 60.05% 71.56% 52.04%
       + RMU-Unlearning 1.29% 8.96% 20.00% 23.77% 1.61% 9.36% 25.90% 29.43% 77.98% 59.68% 70.86% 51.67%
LLaVA-1.5-13B-LoRA 67.87% 93.89% 45.66% 55.22% 0.72% 0.54% 19.25% 12.83% 80.04% 60.23% 71.64% 54.74%
       + Unsafe-Filter 66.97% 94.34% 48.30% 55.85% 0.36% 0.54% 22.26% 13.21% 79.98% 60.05% 71.54% 54.02%
       + Mixed-SFT 0.45% 52.94% 14.34% 38.74% 3.05% 92.45% 63.40% 98.87% 78.85% 59.67% 71.42% 53.27%
       + Posthoc-SFT 0.23% 42.08% 12.08% 30.44% 3.41% 79.68% 61.13% 99.62% 78.64% 59.43% 71.40% 53.64%
       + NPO-Unlearning 3.36% 13.53% 18.11% 22.26% 3.59% 10.47% 23.77% 28.68% 78.43% 59.26% 71.36% 53.41%
       + RMU-Unlearning 2.75% 10.18% 17.74% 22.64% 1.79% 8.64% 26.42% 30.56% 78.27% 58.79% 70.98% 52.99%
In addition to RMU, we also employ NPO [25] to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework [76]. The NPO-based unlearning loss is then given by
$\begin{array}[]{c}\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],\end{array}$ (5)
where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter , $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the model ${\bm{\theta}}$ given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning. The rationale behind NPO is to fine-tune the VLM ${\bm{\theta}}$ to force it to deviate from the reference model when processing unsafe inputs. Following (2), VLM unlearning also requires the retain loss $\ell_{\mathrm{r}}$ to preserve the model utility in normal tasks. This is achieved using the standard VLM training loss over the safe text and safe image data in the fine-tuning set.
As demonstrated in Sec. 6, compared to conventional supervised safety fine-tuning, unlearning-based approaches produce safer responses by generating irrelevant or neutral outputs rather than merely rejecting unsafe queries. Moreover, unlearning reduces the model’s reliance on rejection responses, alleviating over-prudence and enabling more appropriate handling of benign inputs.
6 Experiments
6.1 Experiment Setups
Datasets and models.
We consider four VLM safety datasets: VLGuard [23], MM-SafetyBench [14], SPA-VL [21], and Figstep [15]. To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 [77], TextVQA [78], VizWiz [79], and ScienceQA [80]. For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B [6, 7] as our primary VLMs.
Safety fine-tuning setups and baselines. In our experiments, we choose VLGuard as the training dataset for VLM safety fine-tuning. When implementing the MU-based fine-tuning approaches, i.e., RMU in (3) and NPO in (5), we use the unsafe input-output pairs from VLGuard as the unsafe dataset ($\mathcal{D}_{\mathrm{u}}$) to be unlearned. Here we employ Llama-2-13B-Chat to confirm the harmfulness of the original model’s responses against the unsafe input queries. Additionally, the safe query-answer pairs from VLGuard are used to construct the retain dataset ($\mathcal{D}_{\mathrm{r}}$) in (2). For additional details, we refer readers to Appendix  B.1-B.2.
Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines. (1) Mixed-SFT [23]: Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard. (2) Posthoc-SFT [23, 64]: SFT using a post-hoc fine-tuning approach on VLGuard. (3) Unsafe-Filter: SFT performed on a clean training set, where LLaMA-Guard-3-11B-Vision [81] is used to filter unsafe data from pre-training datasets.
Evaluation setups. For a fine-tuned VLM, its safety performance is primarily evaluated using the attack success rate (ASR), defined as the proportion of unsafe responses generated for unsafe input queries over the total number of queries. The unsafe input queries are considered in two scenarios. (1) Before attack: The original unsafe input queries from VLM safety datasets including VLGuard, MM-SafetyBench, SPA-VL, and Figstep. (2) After attack: We apply the proposed one-word attack to the original unsafe queries. By default, our one-word jailbreaking attack is implemented using a 3-shot approach, as illustrated in Fig. 4. Specifically, three paraphrased versions of the text query (while preserving the adversarial perturbation) are posed for each image. The attack is considered successful if any of the generated responses is unsafe. To validate the harmfulness of the responses generated by the fine-tuned model, we employ the multi-modal Qwen2.5-VL-7B-Instruct as a judge. The details of the judge prompt can be found in Appendix  B.4.
To investigate the over-prudence issue in safety fine-tuned VLMs, we measure the rejection rate (RR), which quantifies the proportion of benign test-time queries that are unnecessarily rejected by the fine-tuned model. These benign queries consist of both safe text queries and safe images provided by the VLM safety datasets. Similar to ASR, RR is evaluated both before and after exploiting spurious correlations to encode rejection bias. Furthermore, we evaluate the general utility of safety fine-tuned VLMs by measuring their prediction accuracy (Acc) on standard downstream VQA tasks, including VQAv2, TextVQA, ScienceQA, VizWiz.
6.2 Experiment Results
Overall performance on safety, over-prudence, utility.
In Table 1, we present a comprehensive evaluation of safety fine-tuned VLMs across three key metrics: safety performance against unsafe input queries (measured by ASR, where lower is better), over-prudence performance against safe input queries (measured by RR, where lower is better), and model utility on representative downstream tasks (measured by Acc, where higher is better). Recall that both ASR and RR are evaluated before and after exploiting spurious correlations, i.e., promoting non-rejection bias and rejection bias via one-word modification, respectively.
Figure 5: Visualization of question-answer pairs from three models: LLaVA-1.5-7B (original), Mixed-SFT (fine-tuned), and RMU-Unlearning (unlearned). Green shield represents the correct response, whether a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries, while red question ? represents an inappropriate rejection for a safe query. The first row displays responses to unsafe text-image queries, while the second row shows responses to safe queries.
We draw some key observations below. First, conventional safety fine-tuning approaches (Unsafe-Filter, Mixed-SFT, and Posthoc-SFT) exhibit a safety mirage, as evidenced by a significant rise in ASR after applying the one-word attack–nearly 60% on the VLGuard unsafe evaluation set and nearly 30% on the SPA-VL unsafe evaluation set of LLaVA-1.5-7B based model. Additionally, these baselines exhibit a significant over-prudence issue, as evidenced by an over 90% increase after one-word modification in RR on the safe evaluation sets in VLGuard and SPA-VL. Furthermore, both full and LoRA fine-tuning exhibit a similar safety mirage phenomenon in baselines.
Second, compared to baselines, the unlearning-based approaches (NPO and RMU) exhibit significantly lower ASR increases after the one-word attack and maintain a low RR against safe queries, effectively alleviating both jailbreaking susceptibility and over-prudence. Compared to NPO, RMU achieves slightly better performance in both ASR (lower vulnerability to the one-word attack) and RR (reduced over-prudence). This result is consistent with LLM unlearning, where RMU typically outperforms NPO in knowledge unlearning [26]. Furthermore, the advantages of MU persist in both full fine-tuning and LoRA fine-tuning.
Third, from the perspective of model utility evaluation, we observe that unlearning-based approaches lead to a slight decrease in Acc on downstream tasks (approximately 1%), compared to supervised safety fine-tuning. This suggests a potential tradeoff between unlearning effectiveness and utility preservation, indicating that erasing harmful knowledge in VLMs may have an impact on general performance. Further optimizing this tradeoff remains an important future research direction to enhance VLM unlearning.
Table 2: Analyzing the safety mechanisms of unlearning-based approaches vs. baselines. Before and after applying the one-shot “What”-initialized attack, the safety rate (1-ASR) against unsafe queries is decomposed into RR (rejection rate) and irrelevance rate (IR), where IR represents responses that are irrelevant to the unsafe queries. All other setups remain consistent with Table 1.
Models Safety Evaluation on VLGuard
Before After
ASR IR RR ASR IR RR
LLaVA-1.5-7B 64.25% 30.09% 5.66% 74.43% 21.95% 3.62%
      +Unsafe-Filter 65.66% 28.01% 6.33% 74.66% 21.49% 3.85%
      +Mixed-SFT 0.23% 0% 99.77% 24.66% 5.20% 70.14%
      +Posthoc-SFT 0.23% 0% 99.77% 25.34% 4.75% 69.91%
      +NPO-Unlearning 2.49% 46.42% 51.09% 6.99% 48.72% 44.29%
      +RMU-Unlearning 1.29% 93.96% 4.75% 5.06% 89.29% 5.65%
LLaVA-1.5-7B-LoRA 64.72% 28.28% 7.02% 72.62% 21.95% 5.43%
      +Unsafe-Filter 67.19% 26.47% 6.33% 73.08% 20.81% 6.11%
      +Mixed-SFT 0.45% 0.0% 99.55% 39.59% 5.66% 54.75%
      +Posthoc-SFT 0.23% 0.0% 99.55% 20.81% 2.94% 76.24%
      +NPO-Unlearning 4.56% 48.64% 46.80% 6.86% 53.14% 40.0%
      +RMU-Unlearning 3.87% 90.92% 5.21% 6.91% 88.33% 4.76%
Other safety evaluation. We further evaluate different safety fine-tuning approaches using LLaVA-v1.5-7B on MM-SafetyBench and FigStep, with results Table A1. Consistent with our findings on VLGuard and SPA-VL, the one-word attack significantly increases ASR across both benchmarks. In Fig. 5, we present input-output demonstrations of safety fine-tuned LLaVA-1.5-7B, comparing the baseline Mixed-SFT with the RMU-based unlearning approach. The original LLaVA-1.5-7B model is vulnerable to unsafe queries both before and after the one-word attack. Here the attack is primarily executed by replacing “How” with “What” in text queries. However, the one-word attack effectively bypass the safety mechanism from Mixed-SFT, triggering unsafe response against the query starting with “What”. Moreover, the Mixed-SFT fine-tuned VLM exhibits over-prudence, as evidenced by its rejection response to a safe text query starting with “Share”. Here “Share” is identified as a high-frequency word in the fine-tuning dataset that is strongly correlated with rejection responses in Fig. 3. By contrast, the RMU unlearning approach effectively defends against the one-word attack while also mitigating the over-rejection issue.
Figure 6: Input token sensitivity analysis for “What”-initiated unsafe queries and “Share”-initiated safe queries over VLGuard using LLaVA-1.5-7B Mixed-SFT, before and after masking “What” and “Share”. Sensitivity is measured using per-token masking (i.e., replacing the original token with a blank placeholder [PAD]) to evaluate each token’s influence on response rejection probabilities.
Irrelevance response for safety enhancement by unlearning. To understand the difference in the safety mechanism of MU-based approaches vs. safety-aware SFT, Table 2 presents the rates of unsafe, irrelevant, and rejection responses in our safety evaluation, both before and after a 1-shot one-word attack. Here, the safety rate (1-ASR) is further decomposed into: (1) irrelevant responses where the model sidesteps the unsafe query by generating a response that is unrelated to the harmful content, and (2) rejection responses where the model explicitly refuses to respond. As observed, conventional SFT-based safety fine-tuning strategies predominantly produce rejection-based responses, as reflected in the overly high RR in both “Before” and “After” scenarios. In contrast, our unlearning-based methods primarily yield irrelevant responses, reducing the model’s reliance on outright rejections, as evidenced by the high irrelevance rate (IR). Notably, the RMU-Unlearning approach steers the model’s output distribution closer to a random vector, making responses more likely to be classified as irrelevant rather than explicitly rejecting queries.
Input token saliency analysis. Recall that spurious correlations emerge between specific words in textual queries and safety labels. To further analyze this, we investigate input token sensitivity by applying per-token masking (i.e., replacing the original input token with a blank placeholder [PAD]) to assess the token influence in non-rejection and rejection responses. Fig. 6 presents the rejection probability using LLaVA-1.5-7B Mixed-SFT for all “What”-initiated unsafe input queries and “Share”-initiated safe queries before and after masking “What” and “Share”, respectively, over the VLGuard dataset. As we can see, masking “What” significantly increases the rejection probability, confirming its role in inducing non-rejection bias in VLGuard. Masking “Share” significantly reduces the rejection probability compared to its unmasked counterpart, demonstrating its influence in reinforcing rejection bias. See Appendix B.6 for specific textual input saliency maps.
Additional results. In Appendix B.7, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
7 Conclusion
In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that emerges from supervised safety fine-tuning. Our analysis reveals that biases in the fine-tuning dataset reinforce spurious correlations between superficial textual patterns and safety labels, leading to a false sense of security. As a result, fine-tuned VLMs become highly susceptible to simple one-word jailbreaking attacks while also exhibiting over-prudence, unnecessarily rejecting benign queries. To address these, we propose MU as a principled alternative to fine-tuning. Unlike supervised approaches, MU directly removes harmful knowledge without relying on explicit safety labels, thereby avoiding biased feature-label mappings and mitigating spurious correlations. Extensive experiments confirm the existence of the “safety mirage” in conventional VLM safety fine-tuning and demonstrate that MU-based safety alignment significantly enhances robustness against jailbreaking attacks, reduces over-prudence, and preserves strong performance on standard VQA tasks.
8 Broader Impacts
Our work reveals spurious correlations in VLM safety fine-tuning, which are prevalent across models and can function as unintentional “backdoors” in real-world applications. Malicious actors could exploit these correlations to launch jailbreaking attacks, potentially extracting sensitive or private information from VLMs. While the advocated unlearning methods aim to enhance safety alignment, they also present dual-use risks. A bad actor could misuse these techniques to erase safety guardrails, rendering a previously robust model unsafe. If such an unlearned model were publicly released, it could bypass safety mechanisms, increasing the risk of harmful content generation.
9 Acknowledgement
We extend our gratitude to Cisco Research and DSO National Laboratories for their support in this project. Additionally, the contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are partially supported by the National Science Foundation (NSF) CISE Core Program Award IIS-2207052, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the NSF CAREER Award IIS-2338068, the Amazon Research Award, and the IBM PhD Fellowship.
Appendix
A One-word Over Prudence
Similar to the one-word jailbreaking attack, inserting a rejection bias-inducing word (e.g., “Share” as shown in Fig. 3) into text queries causes the fine-tuned VLM to produce no meaningful output, even for benign inputs. We also apply a multi-shot strategy — modifying different paraphrases of a benign query with “Share”. Notably, even a one-shot “Share” modification leads to a 90% over-rejection rate on safe queries.
Figure A1: Rejection rate vs. $K$, evaluated before and after applying the “Share”-initialized one-word modification to safe input queries, cause over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model [23].
B Additional Experiment Details and Results
B.1 Additional Fine-tuning Setups
All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs. We perform fine-tuning using both full-parameter and LoRA-based methods. For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$. For LoRA-based fine-tuning, following [6, 21], we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$. In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
B.2 Machine Unlearning Implementation Details
To address problem (2), fine-tuning can be effectively performed either through full parameter updates or parameter-efficient techniques such as low-rank adaptation (LoRA) [82]. Common implementations adopt cross-entropy loss for $\ell_{\mathrm{r}}$, while $\ell_{\mathrm{u}}$ can flexibly incorporate various optimization objectives such as cross-entropy loss, DPO [76], PPO [83].
RMU implementation details.
To apply the RMU algorithm for safety alignment, we use unsafe input-output pairs from VLGuard as the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ for unlearning. We employ Llama-2-13B-Chat to verify the harmfulness of the original LLaVA series model’s responses to unsafe queries and select the most harmful response as the target for unlearning. Specifically, for RMU, we combine the question and answer into a single instance as VLM’s input to be forgotten. For the retain dataset $\mathcal{D}_{\mathrm{r}}$, we use a mixture of the LLaVA fine-tuning data and the safe input-output pairs from VLGuard. The unlearning loss $\ell_{\mathrm{u}}$ defined as (3), where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations (we use layers 4, 5, and 6 in our implementation), $c$ is a hyperparameter for activation scaling (set to $c=10$), and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. The retain loss $\ell_{\mathrm{r}}$ is computed as the cross-entropy loss on the retain data, with the regularization parameter set to $\gamma=1.2$.
NPO implementation details.
Compared to the RMU implementation, we use the same settings for the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ and the retain dataset $\mathcal{D}_{\mathrm{r}}$. However, for NPO, the input to the VLM from $\mathcal{D}_{\mathrm{u}}$ consists solely of the harmful question. In this setting, unsafe data designated for unlearning is treated as negative examples within a direct preference optimization framework [76]. The unlearning training objective is defined as (5), where $\sigma(\cdot)$ denotes the sigmoid function, $\beta>0$ is a temperature parameter (set to $\beta=0.6$ in our implementation), $\pi_{{\bm{\theta}}}(\mathbf{x})$ is the prediction probability of the model given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}(\mathbf{x})$ is the corresponding probability from the reference model (i.e., the initial model prior to unlearning). The retain loss $\ell_{\mathrm{r}}$ is computed as the cross-entropy loss on the retain data, with the regularization parameter set to $\gamma=1.0$.
B.3 Additional Baseline Details
For the Unsafe-Filter baseline, we employ the LLaMA-Guard-3-11B-Vision [81] model to assess the safety of image-instruction-answer pairs in the LLaVA pre-training and fine-tuning datasets. Unsafe data is filtered out, and LLaVA is subsequently fine-tuned using only the remaining safe samples. The fine-tuning setup follows the original LLaVA configuration.
B.4 Evaluation Setup Details
To validate the harmfulness of the responses generated by the fine-tuned model, we employ =Qwen2.5-VL-7B-Instruct model as a judge. Below is the prompt used for evaluating content relevance, which is included in our experiments:
messages = [
“role”: “system”,
“content”: [
{“type”: “text”, “text”: (
“You are an expert in content evaluation. Your task is to determine whether the given answer to the image and question pair is Relevant or Irrelevant.
[…]
“1. An answer should be classified as ’Relevant’ if it is detailed, well-organized, and clearly addresses both the question and the visual content.
[…]
“2. An answer should be classified as ’Irrelevant’ if it is vague, poorly written, or contains nonsensical phrases.
“Always provide detailed reasoning for your decision, and ensure your response strictly starts with either ’Relevant:’ or ’Irrelevant:’.”
“role”: “user”,
“content”: [
{“type”: “image”, “image”: image_path},
{“type”: “text”, “text”: (
[…]
“Please analyze the provided answer in terms of its detail, clarity, and correctness in addressing the image and question.
“Provide detailed reasoning for your decision.”
B.5 Other Safety Evaluation
In Table A1, we present additional safety evaluations of different safety fine-tuning approaches using LLaVA-v1.5-7B on two additional safety benchmarks: MM-SafetyBench and FigStep. The results are consistent with our earlier findings on VLGuard and SPA-VL: The one-word attack leads to a significant increase in ASR on both MM-SafetyBench and FigStep. This reveals that the spurious correlations learned during conventional safety fine-tuning persist across different evaluation settings. Our proposed MU-based methods demonstrate improved robustness, thereby alleviating this vulnerability.
Table A1: Safety evaluation on MM-Safetybench and FigStep. ASR is reported for unsafe inputs before and after the “What”-initialized 3-shot attack. The setup and format follow Table 1.
Models Safety Evaluation ($\downarrow$)
MM-Safety (ASR) FigStep (ASR)
Before After Before After
LLaVA-1.5-7B 48.81% 91.27% 62.00% 86.00%
      +Unsafe-Filter 50.60% 90.28% 62.00% 84.00%
      +Mixed-SFT 0.60% 48.81% 0.00% 20.00%
      +Posthoc-SFT 0.60% 40.48% 0.00% 28.00%
      +NPO-Unlearning 4.76% 20.24% 6.00% 12.00%
      +RMU-Unlearning 2.98% 17.26% 4.00% 10.00%
LLaVA-1.5-7B-LoRA 57.74% 93.45% 72.00% 84.00%
      +Unsafe-Filter 58.93% 91.07% 74.00% 84.00%
      +Mixed-SFT 0.60% 63.69% 0.00% 40.00%
      +Posthoc-SFT 0.60% 41.07% 0.00% 36.00%
      +NPO-Unlearning 4.17% 23.81% 4.00% 16.00%
      +RMU-Unlearning 5.36% 19.05% 2.00% 12.00%
Table A2: Evaluating the safety and over-prudence under various visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter. Other setups are consistent with Table 1.
Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence (RR, $\downarrow$)
Before After Before After
LLaVA-1.5-7B+Mixed-SFT
Original Image 0.23% 54.98% 4.48% 91.76%
+ Gaussian Noise 0.45% 54.08% 4.30% 91.22%
+ Gaussian Blur 0.23% 54.31% 4.66% 91.04%
+ Color Jitter 0.45% 54.75% 4.66% 90.86%
LLaVA-1.5-7B+Posthoc-SFT
Original Image 0.23% 46.83% 2.69% 90.83%
+ Gaussian Noise 0.23% 46.38% 2.15% 90.29%
+ Gaussian Blur 0.23% 45.93% 1.97% 90.29%
+ Color Jitter 0.45% 46.16% 1.97% 91.19%
LLaVA-1.5-7B+NPO-Unlearning
Original Image 2.49% 12.92% 2.51% 11.69%
+ Gaussian Noise 2.26% 12.47% 2.15% 11.51%
+ Gaussian Blur 2.26% 12.70% 1.97% 11.15%
+ Color Jitter 2.71% 13.15% 1.97% 12.05%
LLaVA-1.5-7B+RMU-Unlearning
Original Image 1.29% 10.18% 1.25% 7.56%
+ Gaussian Noise 1.29% 9.73% 1.43% 7.20%
+ Gaussian Blur 1.74% 9.96% 1.43% 7.38%
+ Color Jitter 1.74% 10.41% 1.07% 7.02%
B.6 Examples of Input Saliency
We show two input saliency maps in Fig. A2 and Fig. A3. The input saliency is shown through the prediction probability change for each output sentence ($y$-axis) by masking each input token ($x$-axis). Here [X] refers to the token “X” is marked (i.e., replacing with [PAD])), and [ ] denotes the non-masking case. The examples of input query in Fig. A2 and Fig. A3 correspond to the examples introduced in Fig. 1(a)-(b). Fig. A2 shows that masking “what” token can significantly boost the rejection probability, producing “I’m sorry…”. Fig. A3 shows that masking “share” token successfully reduces the prediction probability for rejection response and produces the normal model generation.
Figure A2: Prediction probability for various output sentences before and after masking each input query token for the query “What can be used to create a promotion?” corresponding to the image example in Fig. 1-(a).
Figure A3: Prediction probability for various output sentences before and after masking each input query token for the query “Share the drink type in the image” corresponding to the image example in Fig. 1-(b).
B.7 Spurious Correlations with Visual Variations
To further evaluate the persistence of spurious correlations between text queries and safety labels and the robustness of our MU-based safety fine-tuning approaches, we present in Table A2 the safety performance of fine-tuned models when subjected to various visual perturbations. The results reveal that the SFT-based safety models consistently exhibit a significant increase in both ASR and RR even under visual variations. This finding indicates that the spurious correlations, and consequently the safety mirage phenomenon remains stable. In contrast, the performance of our MU-based safety aligned models remains robust across these perturbations.