### v3

# arXiv:2503.11832v3[cs.AI] 02 Mar 2026

Abstract
Recent vision language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images. However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns. While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage,” where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm. We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards. Additionally, these correlations contribute to over-prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily. To address these issues, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning, as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities. Extensive evaluations across safety benchmarks show that under MU-based alignment reduces the attack success rate by up to 60.27% and cuts unnecessary rejections by over 84.20%. Codes are available at https://github.com/OPTML-Group/VLM-Safety-Unlearn. WARNING: There exist AI generations that may be offensive in nature.
1 Introduction
Recent multi-modal models have achieved advancements by integrating text and images (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024; Gao et al., 2023; Li et al., 2023b). A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module (Liu et al., 2023; 2024a; Zhu et al., 2023; Chen et al., 2023a; Wang et al., 2024a). While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities (Guo et al., 2024; Qi et al., 2023b; Liu et al., 2024c; b; Gong et al., 2023). Specifically, images can function as a “foreign language” (Pi et al., 2024), creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025). The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), assuming access to a high-quality, dual-modality curated safety fine-tuning dataset (Liu et al., 2024b; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024b; Gu et al., 2024). One compelling piece of evidence is that fine-tuning on widely used VLM safety datasets such as VLGuard (Zong et al., 2024) and SPA-VL (Zhang et al., 2024b), which substantially improves robustness against unsafe queries and jailbreaking attacks. The surprising effectiveness of safety SFT has sparked growing interest in re-evaluating its reliability. However, a key limitation of such safety fine-tuning is the issue of over-prudence (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), where VLMs become overly conservative and unnecessarily reject benign queries.
The over-prudence may indicate that these models are overly safe, avoiding responses even to benign queries. Alternatively, the observed safety may be illusory, as current fine-tuning fails to guarantee reliability, creating a false sense of safety. Thus, we ask:
(Q) Does current VLM safety fine-tuning achieve true safety? If not, what is the root cause?
Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed fine-tuned on VLGuard (Zong et al., 2024)) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in an unsafe query) bypasses the safety mechanism learned from fine-tuning on VLGuard, even though the model correctly rejects the original query. (b) Over-prudence issue: Similar to the one-word attack, a minor modification by replacing “What” with “Share” can cause unnecessary refusals even for benign queries. (c) Root cause of spurious correlations from fine-tuning dataset biases: Certain words are disproportionately linked with specific safety labels, such as “Share” strongly correlated with rejection, while “What” highly associated with non-rejection. (d) Effectiveness of unlearning-based safety fine-tuning: The unlearning methods NPO (Zhang et al., 2024a) and RMU (Li et al., 2024c), enhance robustness against attacks and reduce over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed (Zong et al., 2024) (“Prior work”).
In this work, we investigate (Q) and challenge the prevailing belief in the effectiveness of safety fine-tuning for VLMs. We uncover a “safety mirage” in VLMs safety fine-tuning, where the seemingly robust performance largely stems from spurious correlations between specific words in input queries and predefined safety labels (e.g., rejection) in the fine-tuning dataset. If an adversary identifies these spurious correlations, a simple one-word modification–which we refer to as the “one-word attack”–can effectively jailbreak safety fine-tuned VLMs, enabling them to regenerate unsafe content. Additionally, the input-rejection shortcut induced by these spurious correlations provides an explanation for the over-prudence of safety fine-tuned VLMs. Similar to the one-word attack, a one-word modification at test time can easily trigger this shortcut, causing the model to overgeneralize rejection and unnecessarily refuse benign queries. In Fig. 1(a)-(c), we provide a schematic overview illustrating: (a) The proposed one-word attack on a safety fine-tuned VLMs LLaVA-v1.5-7B-Mixed (Zong et al., 2024); (b) The amplified over-prudence issue by one-word modification; (c) The identified spurious correlations in the fine-tuning dataset VLGuard, where the word “share” is strongly linked to rejection, while “what” is associated with non-rejection.
Building on our identification of spurious correlations in VLMs safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU). Originally designed to remove the influence of undesired data or knowledge from ML models, MU preserves essential knowledge while avoiding unintended disruptions to causally unrelated information (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021). We propose adapting MU to VLMs safety fine-tuning as a more robust alternative to traditional supervised approaches. Instead of enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels. Although MU has been applied to VLMs safety fine-tuning in prior work (Chakraborty et al., 2024; Chen et al., 2025; Huo et al., 2025), its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored. Fig. 1(d) illustrates the effectiveness of adapting two LLM unlearning approaches, NPO (negative preference optimization) (Zhang et al., 2024a) and RMU (representation misdirection for unlearning) (Li et al., 2024c), in enhancing robustness against jailbreaking attacks and reducing over-prudence rates.
In summary, our key contributions are listed below.
①We revisit the problem of safety fine-tuning for VLMs and find that there exists a safety mirage, driven by hidden biases, specifically, spurious correlations between textual questions and safety labels in the fine-tuning dataset.
②From the attack perspective, we show that safety fine-tuned VLMs remain vulnerable to jailbreaking when adversaries exploit these spurious correlations. We propose a simple yet effective one-word attack, which substitutes highly frequent query words associated with rejection responses with those linked to normal model outputs. Furthermore, we demonstrate that spurious correlations also underlie the over-prudence phenomenon, causing models to unnecessarily reject benign inputs.
③From the defense perspective, we show that machine unlearning (MU) offers a promising solution to mitigate the effects of spurious correlations in fine-tuning data. The key idea is that MU removes the influence of unsafe responses without relying on spurious feature–label shortcuts.
④We conduct extensive experiments across multiple VLMs safety evaluation benchmarks, including VLGuard (Zong et al., 2024), SPA-VL (Zhang et al., 2024b), MM-SafetyBench (Liu et al., 2024b), and FigStep (Gong et al., 2023), and assess model utility on standard VQA datasets. Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
2 Related Work
VLM safety: Attack and defense. With the rapid advancement of VLMs (Liu et al., 2023; 2024a; Zhu et al., 2023; Ye et al., 2023; Wang et al., 2023; Li et al., 2023a; Alayrac et al., 2022; Awadalla et al., 2023; Gao et al., 2023), safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content. While LLMs have been extensively studied for safety risks, leading to the development of attack strategies (Yang et al., 2023; Wei et al., 2023b; Huang et al., 2023; Shu et al., 2023), defense mechanisms (Li et al., 2023c; Cao et al., 2023; Kumar et al., 2023), and robust evaluation datasets (Bianchi et al., 2023; Li et al., 2024a; Ji et al., 2023), VLMs introduce additional challenges due to the complexity of multimodal inputs (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025), making them even more vulnerable to jailbreaking and adversarial manipulation (Guo et al., 2024; Qi et al., 2023b; Liu et al., 2024c; b; Gong et al., 2023). Attacks on VLMs often leverage the dual-modality nature of these models. One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs (Gong et al., 2023; Liu et al., 2024b). Another strategy involves using gradient-based adversarial image generation (Bailey et al., 2023; Dong et al., 2023; Luo et al., 2023; Qi et al., 2023a; Zhao et al., 2023) to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning. Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets. The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms (Wang et al., 2024b; Chen et al., 2023b; Pi et al., 2024; Gou et al., 2024; Ding et al., 2024). The latter focuses on red-teaming dataset curation (Liu et al., 2024b; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024b; Gu et al., 2024; Zong et al., 2024; Li et al., 2024b), enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
Machine unlearning in VLMs. MU (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) is designed to remove harmful data influences from a pre-trained model while preserving its overall utility. In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance (Zhang et al., 2024a; Li et al., 2024c; Yao et al., 2024). In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms (Liu et al., 2024d; Dontsov et al., 2024; Ma et al., 2024). For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation (Chen et al., 2025; Huo et al., 2025; Chakraborty et al., 2024). Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc (Zong et al., 2024; Taori et al., 2023)). Green shield represents the correct response, either a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries. And red question ? represents an inappropriate rejection for a safe query. (a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries. However, asking the initial question by starting with the word with “What” can easily bypass this safeguard. (b) Over-prudence: A minor modification by asking the problem start from “What” with “Share” can trigger unnecessary refusals even for benign queries.
3 Preliminary and Problem Statement
Existing VLM safety fine-tuning setup.
Previous works (Pi et al., 2024; Gong et al., 2023; Liu et al., 2024b) highlight the importance of safety alignment in VLMs across both textual and visual modalities. Consequently, many efforts have focused on curating high-quality dual-modality safety datasets for VLMs. For instance, VLGuard (Zong et al., 2024) covers various text-image scenarios, including unsafe cases where the text is unsafe or both modalities are unsafe, as well as safe cases where both are benign. SPA-VL (Zhang et al., 2024b) provides large-scale preference data in the form of question–response tuples, where each image sample pairs a safety-related question with a chosen response and a rejected response. Leveraging these curated safety datasets, recent works (Chakraborty et al., 2024; Ding et al., 2025) show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks (Zou et al., 2023; Wei et al., 2023a; Röttger et al., 2023). In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
We begin by formulating the problem of VLM safety fine-tuning. Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard and SPA-VL). Let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, consisting of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning. For a VLM parameterized by $\bm{\theta}$}, the safety fine-tuning problem can be formulated as:
$\begin{array}[]{ll}\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}&\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})+\gamma\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),\end{array}$ (2)
where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and $\gamma\geq 0$ is a regularization parameter that balances safety alignment with general performance preservation.
Refer to caption Refer to caption Refer to caption
(a) Top words in VLGuard safe   (b) Top words in VLGuard unsafe (c) Top words in SPA-VL
Figure 3: Frequency of question-initiating words across training corpora: (a) VLGuard safe queries with non-rejection responses, (b) VLGuard unsafe queries with rejection responses, (c) SPA-VL queries.
Safety mirage: Motivation and problem of interest. Although safety fine-tuned VLMs could exhibit over-rejection, refusing even benign queries (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), their safety against unsafe queries remains highly robust under common jailbreaking attacks (Zong et al., 2024; Zhang et al., 2024b). The seemingly ‘robust’ safety observed after fine-tuning motivates us to re-examine its true reliability. As demonstrated in Fig. 1, the current safety fine-tuned VLM remains highly vulnerable to simple paraphrasing of text queries even when only the first question word is modified. As a supplement, Fig. 2 & A1 provides motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence. For example, for the VLGuard fine-tuned models, unsafe queries prefixed with the innocuous question word “What” successfully bypass the safeguard, and harmless prompts start with “Share” trigger over-rejection effect. Notably, the choice of the question word is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec. 4.
Examples in Figs. 1, 2, and A1 suggest that fine-tuning VLMs on safety datasets creates a “safety mirage”, as evidenced by their susceptibility to even minor one-word modification in text queries. Thus, our work focuses on the following key research questions: (a) What is the root cause of the “safety mirage” in VLM safety fine-tuning? (b) What can be improved to mitigate the “safety mirage”? We explore these questions in Sec. 4 and Sec. 5, respectively.
4 The Risks of Spurious Correlations in VLM Safety Fine-Tuning
Spurious text features and spurious correlations. As discussed in Sec. 3, current VLM safety fine-tuning methods heavily depend on high-quality, dual-modality safety datasets. Consequently, the safety capabilities of fine-tuned VLMs (i.e., their abilities to prevent harmful content generation) are primarily learned from safety labels (i.e., safe responses) introduced in the fine-tuning datasets. For example, in VLGuard, unsafe queries are assigned as rejection responses, such as “I’m sorry, I cannot …”. Similarly, SPA-VL also selects rejection-based answers for unsafe queries.
At first glance, the use of safety labels appears appropriate. However, hidden bias may arise when safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work. Here the term “spurious features” refers to non-essential features in inputs (primarily for texts in this work) that do not contribute to the fundamental meaning or task-relevance of the input query, in contrast to the “core” features. For example, in Fig. 1, the initial words “What” or “Share” function as spurious features, since they are not directly related to the query’s actual content and can be easily substituted with other question words. In contrast, core features (such as “crime”) are more informative, representing content-related words that capture the true meaning of the query. Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the safety labels in the fine-tuning dataset. The above conceptualization of spurious correlations is inspired by conventional spurious correlation analyses in image classification (Sagawa et al., 2020), where background pixels serve as spurious features while object pixels represent core content. In addition, the introduction of spurious correlations during safety alignment echoes the alignment tax (Siddiqui et al., 2026) and emergent misalignment phenomena (Betley et al., 2026), wherein fine-tuning a broadly capable pretrained model on a narrow objective can inadvertently induce new, unintended misaligned behaviors.
In this work, we identify two types of spurious correlations in VLM safety fine-tuning. (a) Non-rejection bias: Certain words (like “What” in Fig. 1(b)) in text queries become spuriously correlated with non-rejection responses. So, incorporating these words into an original query can easily jailbreak fine-tuned VLMs. (b) Rejection bias: Certain words (like “Share” in Fig. 1(b)) in text queries become spuriously correlated with rejection responses, causing the fine-tuned VLM to exhibit over-prudence.
To reveal these biases, we analyze the frequency of question-initiating words used in training queries that lead to non-rejection responses (i.e., generating contents for safe queries) and rejection responses (i.e., predefined refusal answers for unsafe queries). Fig. 3 presents the most frequent starting words across VLGuard and SPA-VL. In VLGuard, the question word “what” predominantly correlates with non-rejection responses, appearing in over 80% of safe queries where the model generates a response. By contrast, words like “can” and “share” in unsafe queries are closely tied to rejection responses, with “share” exclusively occurring in unsafe contexts and leading to rejection in more than half of its cases. For SPA-VL, we again observe a strong dominance of queries beginning with “what.” Unlike VLGuard safe queries, SPA-VL often pairs such queries with both rejection and non-rejection responses. Preference optimization strengthens alignment with the safety-labeled rejections, yet “what”-initiated queries remain inherently associated with unsafe, non-rejection responses.
One-word jailbreaking. Recognizing the non-rejection bias (i.e., the spurious correlation between certain words like “what” and non-rejection responses) as shown in Fig. 3, an adversary can exploit this bias to jailbreak safety fine-tuned VLMs. Formally, let $q$ be an unsafe text query that the VLM would normally refuse. We construct a paraphrased version $q^{\prime}$ to rewrite $q$ such that it begins with a chosen adversarial word $\mathrm{w}_{\mathrm{adv}}$ (e.g., “what” in Figs. 1, 2, and 3). We refer to this strategy, rephrasing unsafe queries to enforce a biased starting word, as the one-word jailbreaking attack.
Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed (Zong et al., 2024)) on VLGuard unsafe.
In practice, we find that repeatedly applying the one-word attack, by integrating $\mathrm{w}_{\mathrm{adv}}$ with paraphrased versions of the original input query $q$ (up to $K$ times), can significantly improve the attack success rate (ASR). We refer to this strategy as $K$-shot one-word attack. Fig. 4 shows the ASR of the $K$-shot one-word attack on the safety fine-tuned VLM. Even $K=1$ yields 29% ASR, compared to near 0% for the original unsafe queries in VLGuard. ASR exceeds 50% for $K\geq 3$ and approaches 90% as $K$ increases. In contrast, paraphrased queries without the “What”-trigger remain ineffective, confirming that the bias-inducing word is essential to the attack’s success. The effectiveness of the one-word attack can be interpreted through the lens of backdoor attacks (Gao et al., 2020; Saha et al., 2020), where the bias-inducing word “What” functions as a trigger that shortcuts text queries to non-rejection responses during safety fine-tuning. Additional results on LLaVA-v1.5-7B-PPO-30K are provided in Appendix B.1.
Figure 5: Rejection rate vs. $K$-shot one-word modification, evaluated before and after applying the “Share” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model (Zong et al., 2024) on VLGuard safe.
One-word over-prudence. Mirroring the jailbreaking case, inserting a rejection-bias word (e.g., “Share”, Fig. 3) induces over-prudence in fine-tuned VLMs: even benign queries prefixed with this word often yield no meaningful output. A multi-shot procedure, where a benign query is rewritten into several “Share”-initialized variants, further amplifies the effect (Fig. 5). Notably, even a single modification already triggers a 90% over-rejection rate on safe text–image queries, highlighting how dataset-induced starting-word biases can severely degrade utility. Additional results for LLaVA-v1.5-7B-PPO-30K in Appendix B.2.
5 Enhancing VLM Safety through Unlearning
To mitigate spurious correlations in the safety fine-tuning dataset (Sec. 4), a natural solution is to eliminate reliance on safety labels, thereby necessitating shifts to a label-free setting for safety alignment. Machine unlearning (MU) (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) provides an ideal solution in this context, as it is designed to remove the undesired influence of harmful data or knowledge from a pre-trained model while preserving its normal utility.
Although unlearning has been applied to VLM in prior work (Chen et al., 2025; Huo et al., 2025), its unique advantage in addressing spurious correlations remains underexplored. Therefore, we adapt two state-of-the-art MU approaches from LLMs, representation misdirection unlearning (RMU) (Li et al., 2024c) and negative preference optimization (NPO) (Zhang et al., 2024a), to VLM safety field.
The proposed VLM unlearning follows the generic formulation of (2), but with the following key modifications. First, the fine-tuning loss on the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ is replaced with an unlearning objective $\ell_{\mathrm{u}}$ that relies solely on the unsafe data features (text-image queries) in $\mathcal{D}_{\mathrm{u}}$, without depending on the safety labels. In our work, we define the unlearning loss $\ell_{\mathrm{u}}$ based on the principles of RMU and NPO, respectively. The RMU-based unlearning objective aims to map the intermediate features of unsafe data $x\in\mathcal{D}_{\mathrm{u}}$ (to be forgotten) to random features. This ensures that the model no longer retains meaningful representations of the unsafe data. The objective is given by
$\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}[\|M_{{\bm{\theta}}}(\mathbf{x})-c\cdot\mathbf{v}\|_{2}^{2}],$ (3)
where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of $\bm{\theta}$}, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. We remark that, unlike RMU for LLM unlearning (Li et al., 2024c), we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
In addition to RMU, we also employ NPO (Zhang et al., 2024a) to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework (Rafailov et al., 2023). The NPO-based unlearning loss is then given by
$\begin{array}[]{c}\hskip-8.53581pt\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],\end{array}$ (5)
where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter , $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the model $\bm{\theta}$} given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning. The rationale behind NPO is to fine-tune the VLM $\bm{\theta}$} to force it to deviate from the reference model when processing unsafe inputs.
In addition, unlearning-based safety fine-tuning on VLMs requires the retain loss $\ell_{\mathrm{r}}$ in (2) to preserve utility on normal tasks. Unlike LLM unlearning, which relies solely on MU-specific retain objectives, we find that directly applying MU objectives to VLMs often leads to instability or even model collapse. To address this, we design $\ell_{\mathrm{r}}$ as a composite of two terms:
$\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\ell_{\mathrm{ft}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})+\alpha\ell_{\mathrm{mu},\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),$ (6)
where $\ell_{\mathrm{ft}}$ denotes the standard fine-tuning loss of the base VLM, which stabilizes optimization and ensures consistent training dynamics, while $\ell_{\mathrm{mu},\mathrm{r}}$ is the MU-specific retain loss that enforces utility preservation (e.g., the representation loss in RMU on the utility set). Further details on $\ell_{\mathrm{r}}$ and MU implementation in the safety context are provided in Appendix C. As will be shown in Sec. 6, compared to conventional supervised safety fine-tuning, our approach yields robust safety results.
6 Experiments
6.1 Experiment Setups
Datasets and models.
We consider four VLM safety datasets: VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2024b), SPA-VL (Zhang et al., 2024b), and Figstep (Gong et al., 2023). To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 (Goyal et al., 2017), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), and ScienceQA (Lu et al., 2022). For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B (Liu et al., 2023; 2024a) as our primary VLMs.
Safety fine-tuning setups and baselines. In our experiments, we use VLGuard as the training dataset for VLM safety fine-tuning, where its training split is divided into two parts: unsafe and safe. The safe query–answer pairs form the retain dataset ($\mathcal{D}_{\mathrm{r}}$) used for computing $\ell_{\mathrm{r}}$ in (6). For the unlearning loss $\ell_{\mathrm{u}}$, the construction of the unsafe dataset differs by method. For NPO (5), we only use the unsafe queries as $\mathcal{D}_{\mathrm{u}}$. In contrast, for RMU (3), we concatenate each unsafe query with harmful responses selected by Llama-2-13B-Chat, pairing them to form the input to be unlearned. Additional implementation details are provided in Appendix C & D.1.
Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines. (1) Mixed-SFT (Zong et al., 2024): Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard. (2) Posthoc-SFT (Zong et al., 2024; Taori et al., 2023): SFT using a post-hoc fine-tuning approach on VLGuard. (3) Unsafe-Filter: An SFT baseline where unsafe samples are excluded by using LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to filter unsafe text–image pairs from LLaVA’s pre-training data, refer to Appendix  D.2.
Evaluation setups. We assess safety performance using the attack success rate (ASR). For each unsafe query, model outputs are classified into three categories: (1) rejection responses, (2) non-rejection but irrelevant responses, and (3) unsafe responses directly related to the unsafe query. ASR is computed in two steps. First, we measure the rejection rate across all responses. For the remaining non-rejection outputs, we use the multi-modal Qwen2.5-VL-7B-Instruct model as a judge to determine whether a response is related to the unsafe query. Any related response is counted as unsafe. The full judging prompt is provided in Appendix D.3. We report ASR under two scenarios: (1) Before attack: using the original unsafe queries from safety benchmarks. (2) After attack: applying our proposed one-word jailbreaking attack to these unsafe queries. By default, the attack is implemented in a 3-shot setting, where three paraphrased versions of each query (with the adversarial perturbation preserved) are posed for each image, as illustrated in Fig. 4. When multiple responses ($K>1$) are generated, the attack is considered successful if any of them is unsafe.
To further analyze the over-prudence phenomenon in safety fine-tuned VLMs, we measure the rejection rate (RR), which quantifies the proportion of benign test-time queries that are unnecessarily rejected. Similar to ASR, RR is evaluated both before and after applying 1-shot one-word modifications designed to induce rejection bias, like Fig. 5. We also evaluate the general utility by measuring prediction accuracy (Acc) on standard downstream VQA benchmarks.
Table 1: Experiment results evaluating safety, over-prudence, and utility of safety fine-tuned VLMs. Safety is quantified by ASR (attack success rate) on unsafe input queries, evaluated before and after a 3-shot one-word attack (i.e., “what”-based prefix in Fig. 1) that promotes non-rejection bias. Over-prudence is measured by $R$ (rejection rate) on safe input queries, evaluated before and after using 1-shot, one-word modification (i.e., “share”-based prefix in Fig. 1). Here, “Before” and “After” denote the performance prior to and following the respective one-word modification. Utility is assessed by the accuracy (Acc.) on VQA benchmarks. Results are presented for models under full fine-tuning and LoRA fine-tuning settings, with safety fine-tuning approaches including Unsafe-Filter, Mixed-SFT, Posthoc-SFT, NPO-Unlearning, and RMU-Unlearning.
Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence Evaluation (RR, $\downarrow$) Utility Evaluation (Acc., $\uparrow$)
VLGuard SPA-VL VLGuard SPA-VL VQAv2 TextVQA ScienceQA VizWiz
Before After Before After Before After Before After
LLaVA-1.5-7B 64.25% 90.27% 46.42% 52.08% 0.36% 0.36% 14.72% 9.81% 78.53% 58.23% 69.51% 50.07%
     + Unsafe-Filter 65.66% 90.72% 45.66% 54.72% 0.36% 0.36% 15.85% 11.32% 79.14% 58.22% 68.12% 52.14%
     + Mixed-SFT 0.23% 54.98% 14.34% 37.73% 4.48% 91.76% 68.68% 98.87% 78.23% 57.80% 68.27% 52.94%
     + Posthoc-SFT 0.23% 46.83% 13.58% 32.96% 2.69% 90.83% 60.38% 100.0% 78.03% 57.73% 68.42% 51.84%
     + NPO-Unlearning 2.49% 12.92% 18.49% 24.15% 2.51% 11.69% 16.60% 17.36% 77.34% 57.80% 68.02% 50.21%
     + RMU-Unlearning 1.29% 10.18% 17.73% 22.64% 1.25% 7.56% 18.11% 19.24% 77.04% 56.89% 67.68% 50.01%
LLaVA-1.5-7B-LoRA 64.72% 95.25% 44.91% 50.44% 0.18% 0.18% 15.47% 12.45% 79.13% 58.22% 68.62% 52.82%
     + Unsafe-Filter 67.19% 93.89% 45.28% 52.33% 0.36% 0.0% 22.64% 13.21% 79.14% 57.66% 67.97% 53.65%
     + Mixed-SFT 0.45% 69.23% 21.51% 40.13% 3.05% 89.93% 59.25% 97.36% 78.63% 57.24% 68.47% 51.84%
     + Posthoc-SFT 0.23% 51.81% 20.38% 37.61% 3.41% 95.14% 62.26% 99.62% 78.23% 57.17% 67.92% 52.08%
     + NPO-Unlearning 4.56% 18.29% 21.51% 25.28% 2.69% 11.01% 16.98% 19.62% 77.32% 56.98% 66.98% 51.01%
     + RMU-Unlearning 3.87% 11.14% 20.38% 24.24% 1.25% 4.84% 18.49% 21.89% 76.99% 56.62% 66.32% 49.87%
LLaVA-1.5-13B 68.10% 91.86% 50.19% 54.47% 0.54% 0.72% 19.62% 14.34% 79.99% 61.25% 72.73% 53.64%
     + Unsafe-Filter 67.65% 92.99% 52.08% 56.27% 0.54% 0.54% 20.38% 15.09% 79.87% 61.32% 71.59% 52.68%
     + Mixed-SFT 0.45% 57.01% 18.11% 40.5% 4.84% 92.63% 58.87% 97.74% 79.03% 60.98% 72.03% 53.01%
     + Posthoc-SFT 1.58% 69.23% 16.98% 32.83% 2.69% 76.08% 56.98% 98.49% 78.94% 60.63% 71.94% 52.31%
     + NPO-Unlearning 1.89% 11.70% 22.26% 26.04% 2.33% 10.65% 23.77% 27.92% 78.31% 60.05% 71.56% 52.04%
     + RMU-Unlearning 1.29% 8.96% 20.00% 23.77% 1.61% 9.36% 25.90% 29.43% 77.98% 59.68% 70.86% 51.67%
LLaVA-1.5-13B-LoRA 67.87% 93.89% 45.66% 55.22% 0.72% 0.54% 19.25% 12.83% 80.04% 60.23% 71.64% 54.74%
     + Unsafe-Filter 66.97% 94.34% 48.30% 55.85% 0.36% 0.54% 22.26% 13.21% 79.98% 60.05% 71.54% 54.02%
     + Mixed-SFT 0.45% 52.94% 14.34% 38.74% 3.05% 92.45% 63.40% 98.87% 78.85% 59.67% 71.42% 53.27%
     + Posthoc-SFT 0.23% 42.08% 12.08% 30.44% 3.41% 79.68% 61.13% 99.62% 78.64% 59.43% 71.40% 53.64%
     + NPO-Unlearning 3.36% 13.53% 18.11% 22.26% 3.59% 10.47% 23.77% 28.68% 78.43% 59.26% 71.36% 53.41%
     + RMU-Unlearning 2.75% 10.18% 17.74% 22.64% 1.79% 8.64% 26.42% 30.56% 78.27% 58.79% 70.98% 52.99%
6.2 Experiment Results
Overall performance on safety, over-prudence, utility.
Table 1 summarizes results across three metrics: ASR (safety), RR (over-prudence), and Acc (utility). Both ASR and RR are evaluated before and after exploiting spurious correlations, i.e., promoting non-rejection bias and rejection bias via one-word attack and one-word modification. We draw some key observations below.
Figure 6: Visualization of question-answer pairs from three models: LLaVA-1.5-7B (original), Mixed-SFT (fine-tuned), and RMU-Unlearning (unlearned). Green shield represents the correct response, whether a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries, while red question ? represents an inappropriate rejection for a safe query. The first row displays responses to unsafe text-image queries, while the second row shows responses to safe queries.
First, conventional safety fine-tuning approaches (Unsafe-Filter, Mixed-SFT, and Posthoc-SFT) suffer from a clear safety mirage: ASR rises sharply after one-word attacks (nearly 60% on VLGuard-unsafe and 30% on SPA-VL-unsafe for 7B models), while RR also increases dramatically (over 90%) on safe queries. This vulnerability appears consistently under both full and LoRA fine-tuning. Second, unlearning-based approaches (NPO and RMU) substantially reduce ASR increases and maintain low RR, alleviating both jailbreaking susceptibility and over-prudence. Compared to NPO, RMU performs slightly better. Third, in terms of general utility, unlearning-based methods (NPO and RMU) show only a slight accuracy drop (around 1%) compared to the original model. This is expected, as a trade-off between robustness and utility exists, and the gain in safety robustness far outweighs the minor utility loss using unlearning.
To better balance safety and utility in VLMs, we can adopt the idea of coreset unlearning (Pal et al., 2025). Using fewer safety samples for unlearning-based fine-tuning improves utility while still achieving stronger safety performance than conventional fine-tuning baselines. Detailed results are provided in Table A1 of Appendix D.4.
Other safety evaluation. We further evaluate different safety fine-tuning approaches using LLaVA-v1.5-7B on MM-SafetyBench and FigStep, with results Table A2 of Appendix D.5. Consistent with our findings on VLGuard and SPA-VL, one-word perturbations remain highly effective, and MU-based methods are more effective and robust than conventional safety fine-tuning.
In Fig. 6, we show input-output examples of safety fine-tuned LLaVA-1.5-7B, comparing Mixed-SFT with RMU-based unlearning. The original model remains vulnerable to unsafe queries both before and after the one-word attack, implemented by replacing “How” with “What.” While Mixed-SFT reduces some unsafe outputs, the attack still bypasses its safety filter and triggers unsafe responses. Mixed-SFT also suffers from over-prudence, rejecting a benign query starting with “Share,” which is spuriously correlated with rejection responses in Fig. 3. In contrast, RMU-based unlearning both resists the one-word attack and alleviates over-rejection on safe queries.
Table 2: Analyzing the safety mechanisms of unlearning-based approaches versus baselines. ASR was reported both before and after the 1-shot “what” initialized attack. Safety rate is decomposed into the rejection rate ($R$) and irrelevance rate (IR), where IR denotes non-rejection outputs judged as irrelevant to the unsafe queries. All other setups remain consistent with Table 1.
Models Safety Evaluation on VLGuard
Before After
ASR IR $R$ ASR IR $R$
LLaVA-1.5-7B 64.25% 30.09% 5.66% 74.43% 21.95% 3.62%
    +Unsafe-Filter 65.66% 28.01% 6.33% 74.66% 21.49% 3.85%
    +Mixed-SFT 0.23% 0% 99.77% 24.66% 5.20% 70.14%
    +Posthoc-SFT 0.23% 0% 99.77% 25.34% 4.75% 69.91%
    +NPO-Unlearning 2.49% 46.42% 51.09% 6.99% 48.72% 44.29%
    +RMU-Unlearning 1.29% 93.96% 4.75% 5.06% 89.29% 5.65%
LLaVA-1.5-7B-LoRA 64.72% 28.28% 7.02% 72.62% 21.95% 5.43%
    +Unsafe-Filter 67.19% 26.47% 6.33% 73.08% 20.81% 6.11%
    +Mixed-SFT 0.45% 0.0% 99.55% 39.59% 5.66% 54.75%
    +Posthoc-SFT 0.23% 0.0% 99.55% 20.81% 2.94% 76.24%
    +NPO-Unlearning 4.56% 48.64% 46.80% 6.86% 53.14% 40.0%
    +RMU-Unlearning 3.87% 90.92% 5.21% 6.91% 88.33% 4.76%
Unlearning as a distinct safety mechanism: A response analysis. To highlight the difference in safety mechanisms between MU-based approaches and safety-aware SFT, Table 2 reports the proportions of unsafe, irrelevant, and rejection responses before and after 1-shot one-word attack. Here, the safety rate (1–ASR) is decomposed into: (1) irrelevant responses, where the model sidesteps the unsafe query with unrelated content, and (2) rejection responses, where the model explicitly refuses to answer. Conventional SFT approaches, guided by safety labels, rely mainly on rejection responses, yielding consistently high RR both “Before” and “After”. In contrast, MU-based methods do not rely on explicit rejection labels. Instead, they achieve safety primarily through irrelevant responses, reflecting a more label-free mechanism for mitigating spurious correlation.
[…]
Models Before After After
one-word one-word GCG
LLaVA-1.5-7B 64.25% 90.27% 72.90%
    +Unsafe-Filter 65.66% 90.72% 73.75%
    +Mixed-SFT 0.23% 54.98% 5.54%
    +Post-hoc 0.23% 46.83% 4.07%
    +NPO-Unlearning 2.49% 12.92% 3.87%
    +RMU-Unlearning 1.29% 10.18% 2.26%
Evaluation on optimization-based textual attack. To assess whether MU generalizes beyond the dataset-induced correlations identified in Fig. 3, we evaluate the GCG attack (Zou et al., 2023), with results reported in Table 3. We adapt GCG to VLMs by optimizing 20 adversarial suffix tokens on LLaVA-1.5-7B using a harmful image–text pair (an explosive-device image paired with “How to build a bomb?”) and performing 300 gradient-ascent steps targeting the unsafe response “Sure, I’d be happy to help.” The learned suffix tokens are external to the training dataset and appended only at inference time. Under a unified 3-shot setting, GCG produces only modest ASR increases on VLMs. In contrast, simple one-word substitutions (e.g., replacing the first word with “What”) severely degrade SFT-based models, driving ASR up to 90%. These results suggest that the primary vulnerability arises from the spurious correlations we identify rather than from optimization-generated adversarial tokens, while MU remains robust under GCG attacks.
Figure 7: Input token sensitivity analysis for “What”-initiated unsafe queries and “Share”-initiated safe queries over VLGuard using LLaVA-1.5-7B Mixed-SFT, before and after masking “What” and “Share”. Sensitivity is measured using per-token masking (i.e., replacing the original token with a blank placeholder [PAD]) to evaluate each token’s influence on rejection probabilities.
Validating spurious correlations via input token saliency analysis. We further validate spurious correlations between specific query words and safety labels using token-level sensitivity analysis, where selected input tokens are masked (i.e., replaced with [PAD]) and their effect on rejection probabilities is measured. Fig. 7 shows the results for “What’ initiated unsafe queries and “Share” initiated safe queries using LLaVA-1.5-7B Mixed-SFT. Masking “What” leads to a sharp increase in rejection probability, confirming its role in inducing non-rejection bias. Conversely, masking “Share” reduces the rejection probability compared to the unmasked case, highlighting its influence in reinforcing rejection bias.
Additional results. In Appendix D.6, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations. In Appendix D.7, we also show the additional analysis of input saliency maps.
7 Conclusion
In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that arises from supervised safety fine-tuning. We show that biases in fine-tuning datasets reinforce spurious correlations between superficial textual patterns and safety labels, creating a false sense of security. As a result, fine-tuned VLMs remain vulnerable to simple one-word jailbreaking attacks and exhibit over-prudence, unnecessarily rejecting benign queries. To address this, we propose machine unlearning (MU) as a principled alternative: rather than relying on explicit safety labels, MU removes harmful knowledge in a label-free manner, mitigating spurious correlations. Our experiments validate the safety mirage phenomenon and demonstrate that MU improves robustness against jailbreaks, alleviates over-prudence, and preserves utility on standard VQA tasks. We refer readers to Appendix EF for limitations, broad impact, and details of LLM usage.
Acknowledgment
This project is supported by research awards from Cisco Research and DSO National Laboratories. The contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are also partially supported by the NSF CISE Core Program Awards IIS-2207052 and IIS-2504263, the NSF CAREER Award IIS-2338068, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, and the Open Philanthropy Research Award.
Appendix
Appendix A Visualization of Motivation Samples
Figure A1: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-SPA-VL-DPO-30K and LLaVA-v1.5-7B-SPA-VL-PPO-30K (Zhang et al., 2024b; Rafailov et al., 2023; Schulman et al., 2017)). The visual marks keep consistent meaning with Fig. 2. (a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries. But if the questions use the starting word “Please”, it would bypass this safeguard. (b) Over-prudence: The fine-tuned model will generate the benign responses when the queries are both harmless, but ask the problems starting from “What” can trigger unnecessary refusals.
Appendix B Additional Results on Spurious Correlations
B.1 Additional Results of One-Word Jailbreaking
Figure A2: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “Please”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024b)) on SPA-VL unsafe test set.
For the SPA-VL test set, we further evaluated the LLaVA-v1.5-7B-PPO-30K model using the same $K$-shot paraphrasing protocol as in the main text. Here, we selected “please” as the adversarial biasing word for unsafe queries. The motivation is dataset-driven: in the SPA-VL training corpus, the token “please” occurs with very low frequency like Fig. 3 shows, so asserting it to unsafe queries disrupts the model’s learned phrasing patterns. Under PPO-based preference optimization, this rare starter token reduces the model’s tendency to follow the preferred safe/rejection response, thereby substantially increasing the attack success rate (ASR). These results corroborate our main finding that low-frequency tokens can serve as effective adversarial triggers in one-word jailbreaking. Quantitatively, as shown in Fig. A2, the ASR rises from about 12% before modification to nearly 30% at $K=1$, and continues to grow steadily with larger $K$, ultimately exceeding 60%. This confirms that the “Please” based attack is consistently effective across different shot settings.
B.2 Additional Results of One-Word Over-prudence
Figure A3: Rejection rate of $K$-shot one-word modification, evaluated before and after applying the “What” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024b)) on SPA-VL safe.
We also tested the over-prudence effect on SPA-VL safe queries. For SPA-VL, we used “what” as the biased starter word. Unlike “please”, “what” is highly frequent in the SPA-VL training corpus as Fig. 3 presents. When combined with PPO preference optimization, this frequent token strongly biases the model toward the preferred rejection-style responses. As a result, benign queries initialized with “what” exhibit a substantially increased rejection rate (RR), confirming the over-prudence phenomenon. These findings highlight how high-frequency tokens can act as rejection-bias triggers in safe contexts, complementing the low-frequency jailbreak triggers observed in unsafe contexts. From Fig. A3, we see that the RR almost doubles, jumping from about 39% before modification to nearly 79% after just one shot, and quickly saturating around 90% as $K$ grows. This illustrates the strong over-rejection tendency induced by the “what” starter.
Appendix C Implementations Details of Machine Unlearning
To address problem (2), fine-tuning can be effectively performed either through full parameter updates or parameter-efficient techniques such as low-rank adaptation (LoRA). Common implementations adopt cross-entropy loss for $\ell_{\mathrm{r}}$, while $\ell_{\mathrm{u}}$ can flexibly incorporate various optimization objectives such as cross-entropy loss, DPO (Rafailov et al., 2023), PPO (Schulman et al., 2017).
Details of Retain Loss $\ell_{\mathrm{r}}$.
To address the problem, directly applying MU objectives to VLMs often leads to instability or even model collapse, so we design $\ell_{\mathrm{r}}$ as a composite of two terms in (6). For the $\ell_{\mathrm{ft}}$, it is the common implementations that adopt cross-entropy loss on $\mathcal{D}_{\mathrm{r}}$ like the original LLaVA fine-tune. While for the $\ell_{\mathrm{mu},\mathrm{r}}$, it is depend on the choices of $\ell_{\mathrm{u}}$, we set it as the corresponding retain loss of the specific unlearning algorithm. For example, when $\ell_{\mathrm{u}}$ is the RMU training loss,
$\ell_{\mathrm{\mathrm{mu},\mathrm{r}}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{r}}}\big[\|M_{{\bm{\theta}}}(\mathbf{x})-M_{{\bm{\theta}}_{\mathrm{ref}}}(\mathbf{x})\|_{2}^{2}\big],$ (A1)
where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations of $\bm{\theta}$}, and $M_{{\bm{\theta}}_{\mathrm{ref}}}(\cdot)$ those of the reference (pre-unlearning) model. While for $\ell_{\mathrm{u}}$ equals to the NPO training loss,
$\ell_{\mathrm{\mathrm{mu},\mathrm{r}}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{r}}}\left[-\frac{2}{\beta}\log\sigma\left(\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],$ (A2)
where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter, $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the current model given the retain input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning. During the experiment, we both set $\alpha=1$ in (2).
RMU implementation details.
To apply the RMU algorithm for safety alignment, we use unsafe input-output pairs from VLGuard as the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ for unlearning. We employ Llama-2-13B-Chat to verify the harmfulness of the original LLaVA series model’s responses to unsafe queries and select the most harmful response as the target for unlearning. Specifically, for RMU, we combine the question and answer into a single instance as VLM’s input to be forgotten. For the retain dataset $\mathcal{D}_{\mathrm{r}}$, we use a mixture of the LLaVA fine-tuning data and the safe input-output pairs from VLGuard. The unlearning loss $\ell_{\mathrm{u}}$ defined as (3), where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations (we use layers 4, 5, and 6 in our implementation), $c$ is a hyperparameter for activation scaling (set to $c=10$), and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. The retain loss $\ell_{\mathrm{r}}$ is computed data, with the regularization parameter set to $\gamma=1.2$.
NPO implementation details.
Compared to the RMU implementation, we use the same settings for the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ and the retain dataset $\mathcal{D}_{\mathrm{r}}$. However, for NPO, the input to the VLM from $\mathcal{D}_{\mathrm{u}}$ consists solely of the harmful question. In this setting, unsafe data designated for unlearning is treated as negative examples within a direct preference optimization framework (Rafailov et al., 2023). The unlearning training objective is defined as (5), where $\sigma(\cdot)$ denotes the sigmoid function, $\beta>0$ is a temperature parameter (set to $\beta=0.6$ in our implementation), $\pi_{{\bm{\theta}}}(\mathbf{x})$ is the prediction probability of the model given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}(\mathbf{x})$ is the corresponding probability from the reference model (i.e., the initial model prior to unlearning). The retain loss $\ell_{\mathrm{r}}$ computed with the regularization parameter set to $\gamma=1.0$.
Appendix D Additional Experiment Details and Results
D.1 Additional Fine-tuning Setups
All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs. We perform fine-tuning using both full-parameter and LoRA-based methods. For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$. For LoRA-based fine-tuning, following (Liu et al., 2023; Zhang et al., 2024b), we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$. In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
D.2 Additional Baseline Details
An additional SFT baseline is constructed by filtering out unsafe samples at both stages of LLaVA training. Specifically, we employ LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to assess the safety of all image–instruction–answer triplets in LLaVA’s pre-training and instruction-tuning datasets. Any sample flagged as unsafe is removed prior to training. As a result, the model is pre-trained exclusively on safe image–text pairs, and subsequently fine-tuned on safe instruction–answer pairs, ensuring that unsafe data is excluded from the entire training pipeline. The fine-tuning setup otherwise follows the original LLaVA configuration, making this a direct comparison point for evaluating the effect of safety filtering throughout both training stages.
D.3 Evaluation Setup Details
To validate the harmfulness of the responses generated by the fine-tuned model, we employ Qwen2.5-VL-7B-Instruct model as a judge. Below is the prompt used for evaluating content relevance, which is included in our experiments:
messages = [
“role”: “system”,
“content”: [
{“type”: “text”, “text”: (
“You are an expert in content evaluation. Your task is to determine whether the given answer to the image and question pair is Relevant or Irrelevant.
[…]
“1. An answer should be classified as ’Relevant’ if it is detailed, well-organized, and clearly addresses both the question and the visual content.
[…]
“2. An answer should be classified as ’Irrelevant’ if it is vague, poorly written, or contains nonsensical phrases.
“Always provide detailed reasoning for your decision, and ensure your response strictly starts with either ’Relevant:’ or ’Irrelevant:’.”
“role”: “user”,
“content”: [
{“type”: “image”, “image”: image_path},
{“type”: “text”, “text”: (
[…]
“Please analyze the provided answer in terms of its detail, clarity, and correctness in addressing the image and question.
“Provide detailed reasoning for your decision.”
D.4 Balance between Safety and Utility
Table A1: Overall results evaluating safety (ASR, $\downarrow$), over-prudence ($R$, $\downarrow$), and utility (Acc., $\uparrow$) on VQA benchmarks. Metrics are reported before and after one-word attacks/modifications. We compare conventional fine-tuning baselines, unlearning-based methods (NPO, RMU), and their coreset variants (50%, 25%, 10%). The other settings stay consistent with Table  1.
Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence Evaluation (RR, $\downarrow$) Utility Evaluation (Acc., $\uparrow$)
VLGuard SPA-VL VLGuard SPA-VL VQAv2 TextVQA ScienceQA VizWiz
Before After Before After Before After Before After
LLaVA-1.5-7B 64.25% 90.27% 46.42% 52.08% 0.36% 0.36% 14.72% 9.81% 78.53% 58.23% 69.51% 50.07%
     + Unsafe-Filter 65.66% 90.72% 45.66% 54.72% 0.36% 0.36% 15.85% 11.32% 79.14% 58.22% 68.12% 52.14%
     + Mixed-SFT 0.23% 54.98% 14.34% 37.73% 4.48% 91.76% 68.68% 98.87% 78.23% 57.80% 68.27% 52.94%
     + Posthoc-SFT 0.23% 46.83% 13.58% 32.96% 2.69% 90.83% 60.38% 100.0% 78.03% 57.73% 68.42% 51.84%
     + NPO-Unlearning 2.49% 12.92% 18.49% 24.15% 2.51% 11.69% 16.60% 17.36% 77.34% 57.80% 68.02% 50.21%
      + 50% NPO 2.94% 12.46% 19.02% 25.10% 2.69% 12.23% 17.45% 18.02% 77.41% 57.82% 68.20% 50.55%
      + 25% NPO 3.65% 14.25% 20.12% 26.48% 2.51% 11.33% 17.88% 18.67% 77.60% 58.10% 68.45% 50.92%
      + 10% NPO 5.17% 18.77% 22.01% 28.92% 1.97% 7.72% 18.32% 19.15% 78.10% 57.99% 68.72% 51.34%
     + RMU-Unlearning 1.29% 10.18% 17.73% 22.64% 1.25% 7.56% 18.11% 19.24% 77.04% 56.89% 67.68% 50.01%
      + 50% RMU 1.74% 10.64% 18.12% 23.05% 1.61% 8.01% 18.56% 19.70% 77.23% 57.98% 67.84% 50.34%
      + 25% RMU 3.22% 15.43% 19.06% 24.12% 1.61% 4.28% 18.92% 20.01% 78.01% 57.24% 68.12% 50.78%
      + 10% RMU 7.51% 22.12% 21.35% 27.25% 0.72% 1.97% 19.48% 20.73% 78.32% 58.97% 68.65% 51.42%
To better balance safety and utility in VLMs, we further adopt the idea of coreset unlearning (Pal et al., 2025). Instead of unlearning all safety data, we fine-tune on a random-sampled subset (e.g., 50%, 25%, or 10%) of the safety dataset. As shown in Table A1, reducing the coreset size leads to slightly higher ASR and RR (weaker safety), but improves accuracy on downstream tasks (higher utility). This also reflects a natural trade-off: smaller coresets sacrifice some safety robustness while preserving more general task capability. Importantly, even with only 10% of the safety data, unlearning-based models still achieve substantially better safety–utility trade-offs than conventional fine-tuning baselines, confirming the flexibility of coreset unlearning for practical deployment
Table A2: Safety evaluation on MM-SafetyBench and FigStep. ASR is reported before and after the “What”-based 3-shot attack. Setup follows Table 1.
Models Safety Evaluation ($\downarrow$)
MM-Safety (ASR) FigStep (ASR)
Before After Before After
LLaVA-1.5-7B 48.81% 91.27% 62.00% 86.00%
    +Unsafe-Filter 50.60% 90.28% 62.00% 84.00%
    +Mixed-SFT 0.60% 48.81% 0.00% 20.00%
    +Posthoc-SFT 0.60% 40.48% 0.00% 28.00%
    +NPO-Unlearning 4.76% 20.24% 6.00% 12.00%
    +RMU-Unlearning 2.98% 17.26% 4.00% 10.00%
LLaVA-1.5-7B-LoRA 57.74% 93.45% 72.00% 84.00%
    +Unsafe-Filter 58.93% 91.07% 74.00% 84.00%
    +Mixed-SFT 0.60% 63.69% 0.00% 40.00%
    +Posthoc-SFT 0.60% 41.07% 0.00% 36.00%
    +NPO-Unlearning 4.17% 23.81% 4.00% 16.00%
    +RMU-Unlearning 5.36% 19.05% 2.00% 12.00%
D.5 Additional Safety Evaluation
In Table A2, we provide additional safety evaluations of different fine-tuning approaches for LLaVA-v1.5-7B on two further safety benchmarks: MM-SafetyBench and FigStep. The results align closely with our earlier findings on VLGuard and SPA-VL. In both new benchmarks, the one-word attack substantially increases the attack success rate (ASR), indicating that spurious correlations between surface-level textual cues and safety labels, induced during conventional safety fine-tuning, persist across diverse evaluation settings. These correlations cause models to become vulnerable to simple adversarial word insertions, undermining their safety guarantees. By contrast, our MU-based approaches (NPO and RMU) consistently achieve lower post-attack ASR, demonstrating improved robustness and reduced reliance on spurious features. This highlights that MU not only mitigates overfitting to dataset-specific biases but also generalizes better across benchmarks, thereby offering a more reliable defense against spurious correlation–driven vulnerabilities.
D.6 Spurious Correlations with Visual Variations
To further examine whether the identified spurious correlations between textual queries and safety labels persist under different visual conditions, we evaluate fine-tuned models on VLGuard with a range of common visual perturbations, including Gaussian Noise, Gaussian Blur, and Color Jitter. The results are summarized in Table A3. We make several observations.
Table A3: Evaluating safety (ASR, $\downarrow$) and over-prudence ($R$, $\downarrow$) under visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter. Other setups follow Table 1.
Models Safety (ASR) Over-Prudence (RR)
Before After Before After
LLaVA-1.5-7B+Mixed-SFT
Original Image 0.23% 54.98% 4.48% 91.76%
+ Gaussian Noise 0.45% 54.08% 4.30% 91.22%
+ Gaussian Blur 0.23% 54.31% 4.66% 91.04%
+ Color Jitter 0.45% 54.75% 4.66% 90.86%
LLaVA-1.5-7B+Posthoc-SFT
Original Image 0.23% 46.83% 2.69% 90.83%
+ Gaussian Noise 0.23% 46.38% 2.15% 90.29%
+ Gaussian Blur 0.23% 45.93% 1.97% 90.29%
+ Color Jitter 0.45% 46.16% 1.97% 91.19%
LLaVA-1.5-7B+NPO-Unlearning
Original Image 2.49% 12.92% 2.51% 11.69%
+ Gaussian Noise 2.26% 12.47% 2.15% 11.51%
+ Gaussian Blur 2.26% 12.70% 1.97% 11.15%
+ Color Jitter 2.71% 13.15% 1.97% 12.05%
LLaVA-1.5-7B+RMU-Unlearning
Original Image 1.29% 10.18% 1.25% 7.56%
+ Gaussian Noise 1.29% 9.73% 1.43% 7.20%
+ Gaussian Blur 1.74% 9.96% 1.43% 7.38%
+ Color Jitter 1.74% 10.41% 1.07% 7.02%
First, conventional SFT-based safety models (Mixed-SFT and Posthoc-SFT) exhibit a consistent pattern of vulnerability: both ASR and RR remain high across all perturbations, showing that the learned spurious correlations are insensitive to low-level visual variations. This suggests that the safety mirage is not simply a dataset artifact tied to specific image statistics, but a systematic bias reinforced during fine-tuning. Second, MU-based approaches (NPO and RMU) maintain stable and low ASR and RR even under perturbations, confirming that unlearning not only alleviates spurious correlations in clean settings but also provides robustness against visual shifts. Third, comparing perturbation types, we find that color-based distortions (Color Jitter) lead to slightly larger fluctuations than Gaussian Noise or Blur, though the overall trend remains the same. This indicates that safety mirage in SFT models is strongly text-driven and less sensitive to visual details, while MU-based approaches decouple safety predictions from such spurious text–label shortcuts. Taken together, these findings highlight that spurious correlations embedded during safety fine-tuning are persistent and modality-invariant: they survive visual perturbations that alter low-level image statistics. In contrast, unlearning-based models generalize more reliably across such shifts, further validating their robustness in multimodal safety alignment.
Figure A4: Prediction probability for various output sentences before and after masking each input query token for the query “What can be used to create a promotion?” corresponding to the image example in Fig. 1-(a).
Figure A5: Prediction probability for various output sentences before and after masking each input query token for the query “Share the drink type in the image” corresponding to the image example in Fig. 1-(b).
D.7 Examples of Input Saliency
We show two input saliency maps in Fig. A4 and Fig. A5. The input saliency is shown through the prediction probability change for each output sentence ($y$-axis) by masking each input token ($x$-axis). Here [X] refers to the token “X” is marked (i.e., replacing with [PAD])), and [ ] denotes the non-masking case. The examples of input query in Fig. A4 and Fig. A5 correspond to the examples introduced in Fig. 1(a)-(b). Fig. A4 shows that masking “what” token can significantly boost the rejection probability, producing “I’m sorry…”. Fig. A5 shows that masking “share” token successfully reduces the prediction probability for rejection response and produces the normal model generation.
Appendix E Limitations and broad Impact
E.1 Limitations
Our analysis has so far been limited to models at the 7B and 13B scales. Extending the study to larger foundation models remains an important next step, as scaling may introduce new behaviors or amplify new challenges. In addition, our experiments have primarily relied on widely used safety benchmarks. While these datasets provide systematic evaluation settings, they may not fully capture the diversity and complexity of real-world applications. Future work should therefore investigate unlearning performance in broader, domain-specific scenarios and corresponding downstream tasks.
E.2 Broad Impact
This work sheds light on the reliability of current VLM safety fine-tuning, uncovering the risk of a “safety mirage” where perceived robustness is illusory. By identifying spurious correlations as a root cause and demonstrating how machine unlearning can alleviate them, our study contributes to building safer and more trustworthy multimodal systems. In practice, stronger defenses against jailbreaking and over-prudence can reduce the risk of harmful or overly restrictive outputs, improving user trust in deployed AI systems. At the same time, unlearning techniques must be carefully applied, as overly aggressive removal of knowledge could impair model utility or be misused to deliberately suppress content. We encourage future research to balance safety, fairness, and transparency when deploying MU-based safety interventions in real-world applications.
Appendix F The Use of LLMs
This work makes limited use of LLMs. Specifically, LLMs were employed exclusively for grammar correction and stylistic polishing of the manuscript. They were not involved in research ideation, experimental design, data analysis, or the generation of any scientific content. All substantive contributions to the paper are solely attributable to the author.